OpenStudy (anonymous):

Differentiate the equation x^3 + y^3 = 3 with respect to y.

OpenStudy (anonymous):

\[3x^{2}\frac{dx}{dy} + 3y^{2} = 0\]\[\frac{dx}{dy} = -\frac{y^{2}}{x^{2}}\][Implicit differentiation with y as independent variable and x as dependent variable].